OpenStudy (anonymous):

how do you graph it: x+y=6 x-y=4

OpenStudy (anonymous):

x+y=6 -x = -x y= -x+6

OpenStudy (anonymous):

x-y=4 -x = -x -y= -x +4 -y/-1 = -x/-1 + 4/-1 y= x-4